Download the Profit and Loss example associated w/ this module (3). Using the notes from this weeks class, complete the pro-forma Profit and Loss statement and upload to this link.

PROFIT AND LOSS STATEMENT

Given: Total Seats

100

Total turns 2

ASSUMPTIONS: Average check: $25 food and $10 beverage.

Food cost is 28% of food sales. Beverage cost is 19% of beverage sales, and labor is 29%.

Direct operating expenses are 28% and occupancy cost is 8%.

What are total

weekly

sales?

What is the SALES MIX as a percentage?

What is food cost as a dollar amount? Remember, you must find food sales first.

What is beverage cost as a dollar amount? Remember, you must find beverage sales first.

What is the cost of goods sold as a percentage?

What is the gross profit as a dollar amount?

What is the gross profit as a percentage?

What is labor cost as a dollar amount?

What are direct operating expenses in dollars?

What are occupancy costs in dollars?

What is the profit in dollars?
Answer #1

What are total weekly sales? = 100 seats * Average check (25+10) * 7 days a week = $ 24,500

What is the SALES MIX as a percentage? Food = 25/35 = 71.4% Beverage = 10/35 = 28.6 %

What is food cost as a dollar amount? = 100*25*7*28% = $ 4900

What is beverage cost as a dollar amount? = 100*10*7*19% = $ 1330

What is the cost of goods sold as a percentage? = (4900+1330)/24500 = 25.4 %

What is the gross profit as a dollar amount> = 24500 - (4900+1330) = $ 18270

What is the gross profit as a percentage? = 18270/24500 = 74.6%

What is labor cost as a dollar amount? = 24500*29% = $ 7105

What are direct operating expenses in dollars? = 24500*28% = $ 6860

What are occupancy costs in dollars? = 24500*8% = $ 1960

What is the profit in dollars? = 18270 - 7105 - 6860 - 1960 = $ 2345